### v1

# Characterization of Multipartite non-Gaussian Entanglement Structure

Abstract
Multipartite entanglement is an essential resource for quantum information tasks, but characterizing entanglement structures in continuous variable systems remains challenging, especially in multimode non-Gaussian scenarios. In this work, we introduce a method for detecting multipartite entanglement structures in arbitrary continuous variable states, which also reveals genuine multipartite entanglement. By leveraging the quantum Fisher information, we propose a criterion for characterizing entanglement and offer a systematic approach to identify feasible operators that capture quantum correlations in multimode non-Gaussian states. We demonstrate the effectiveness of our method on over $10^{4}$ randomly generated multimode-entangled quantum states, achieving a high success rate in entanglement detection. Additionally, our method exhibits enhanced robustness against losses by expanding the set of accessible operators. This work provides a general framework for characterizing entanglement structures in diverse continuous variable systems, enabling a number of experimentally relevant applications.
Multipartite entanglement is an essential resource for many quantum computation, sensing and communication protocols [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Its importance is particularly evident in continuous variable (CV) systems, where multimode Gaussian states, such as squeezed states and cluster states, can be deterministically prepared and controlled [11, 12, 13]. Moreover, non-Gaussian features in CV systems are proven to be a resource providing an advantage over classical protocols [14]. These non-Gaussian states play a crucial role for quantum computation [15, 16, 17], sensing [18, 19, 20, 21], and imaging [22, 23]. Strategies to generate and control such states has seen substantial advancements in recent years, including through photon-subtraction [24, 25], spontaneous parametric down-conversion (SPDC) [26, 27], as well as in microwave cavities coupled to superconducting artificial atoms [28, 29].
Despite these advancements, accurately characterizing both Gaussian and non-Gaussian multimode entanglement in CV systems continues to present substantial challenges [14], e.g. due to the complex correlations appearing among higher-order statistical moments of quadrature field operations. Unlike discrete variable (DV) systems, where entanglement structures are identified based on separable partitions and their sizes [30, 31, 32], there is no such comprehensive method for CV systems. Conventional works [33, 34, 35] in CV systems can only distinguish entanglement structures for the one given partition and become ineffective with convex combinations of different partitions. This limitation hinders the understanding of complex structures, such as m-mode genuine multipartite entanglement. Moreover, while quantum correlations in Gaussian states are predominantly analyzed through first and second-order moments [36, 34, 37], non-Gaussian states involve more complex higher-order correlations. In general, these correlations can not be captured by standard criteria based on Gaussian metrics, which typically focus on the first- and second-order correlations alone. Therefore, there is a pressing need to develop a method to characterize entanglement structure in CV systems, filling this significant gap in the field.
Here, we propose a general method to characterize entanglement structure based on quantum metrology tools. Specifically, we systematically determine the operators that can effectively capture complex correlations in multimode systems. Our approach begins with the analytical identification of a convex set of operators, each capable of witnessing specific bi-inseparable entanglement structures across arbitrary quantum states. Since different entanglement structures correlate with distinct convex sets, we then proceed to pinpoint the intersection of these sets. This intersection allows us to identify a relatively optimal operator for witnessing entanglement. Crucially, contrary to previous work which allows for identifying a specific entanglement structure, our method allows for identifying also their convex combinations, which enables the classification of entanglement and detection of genuine multipartite entanglement (GME). To demonstrate the efficiency of our method, we apply it to a broad range of non-Gaussian states, generated through various stellar ranks and nonlinear processes, including over $10^{4}$ randomly produced quantum states. Additionally, we examine the robustness of our method in scenarios involving loss, demonstrating that the inclusion of high-order operators significantly enhances resilience to loss channels. This approach marks the first accessible method for witnessing entanglement structure in arbitrary Gaussian and non-Gaussian states, thereby addressing a significant need in CV systems.
Figure 1: Graph representation of different multipartite entanglement structures. As illustrated, a system with $N=4$ modes can be characterized by 15 different partitions, associated to different distributions of correlations as indicated by the solid circles. Each partition can be graphically represented by a Young diagram, where the number of boxes in each row indicates the largest of correlated modes, while the number of boxes in the first column indicates the number of partitions of the system.
Metrological bound for $\Lambda$-partitions.– We begin discussing the entanglement structure for $N$ parties. A partition $\mathcal{K}=\left\{H_{1},H_{2},\ldots,H_{k}\right\}$ separates the total $N$-partite system into $k$ nonempty, disjoint subsets $H_{l}$, where each subset has a size $N_{l}$ such that $\sum_{l=1}^{k}N_{l}=N$. A state $\hat{\rho}_{\mathcal{K}}$ is $\mathcal{K}$-separable if there exist local quantum states $\hat{\rho}_{H_{l}}^{(\gamma)}$ for each subset and a probability distribution $p_{\gamma}$ such that $\hat{\rho}_{\mathcal{K}}=\sum_{\gamma}p_{\gamma}\hat{\rho}_{H_{1}}^{(\gamma)}\otimes\cdots\otimes\hat{\rho}_{H_{k}}^{(\gamma)}$. A partition $\mathcal{K}$ can be classified according to the number of subsets and the (largest) size of the subsets. This can be illustrated graphically by means of Young diagrams, as illustrated in Fig. 1 for a 4-mode system. We employ the notation $\Lambda$ to refer to convex combinations of partitions associated with identical Young diagrams. And any states that cannot be written as a $\Lambda$-separable state $\hat{\rho}_{\Lambda}=\sum_{\mathcal{K}\in\Lambda}P_{\mathcal{K}}\hat{\rho}_{\mathcal{K}}$ are denoted as $\Lambda$-inseparable states [38].
Using the framework of quantum metrology, it is known that for an arbitrary $\mathcal{K}$-separable state $\hat{\rho}_{\mathcal{K}}$, the quantum Fisher information (QFI) is bounded by the variances as [39, 33]
$F_{Q}[\hat{\rho}_{\mathcal{K}},\hat{A}]\leq 4\mathcal{V}(\hat{\rho}_{\mathcal{K}},\hat{A}).$ (1)
Here, $\hat{A}=\sum_{j=1}^{N}\hat{A}_{j}$ with $\hat{A}_{j}$ represents a local generator in the reduced state $\hat{\rho}_{j}$, $\mathcal{V}(\hat{\rho}_{\mathcal{K}},\hat{A})=\sum_{i=1}^{k}\mathrm{Var}(\hat{\rho}_{H_{i}},\sum_{j\in H_{i}}\hat{A}_{j})$, and $\mathrm{Var}(\hat{\rho},\hat{O})=\langle\hat{O}^{2}\rangle_{\hat{\rho}}-\langle\hat{O}\rangle^{2}_{\hat{\rho}}$ defines the variance. The explicit expression of QFI is given by $F_{Q}[\hat{\rho}_{\mathcal{K}},\hat{A}]=2\sum_{\begin{subarray}{c}k,l\\ p_{k}+p_{l}\neq 0\end{subarray}}\frac{(p_{k}-p_{l})^{2}}{p_{k}+p_{l}}|\langle\Psi_{k}|\hat{A}|\Psi_{l}\rangle|^{2}$, with the spectral decomposition $\hat{\rho}=\sum_{k}p_{k}|\Psi_{k}\rangle\langle\Psi_{k}|$ [40]. Since Eq. (1) is a necessary criterion for separability, its violation implies $\mathcal{K}$-paritite inseparability.
To distinguish different $\Lambda$ partitions, a general metrological bound that is suitable for all mixed $\mathcal{K}$-separable states with $\mathcal{K}\in\Lambda$ needs to be given.
Theorem 1. $\Lambda$-partite inseparability can be confirmed if there is at least one operator $\hat{A}$ that can violate Eq. (1) across all $\mathcal{K}$-partition with $\mathcal{K}\in\Lambda$.
Proof.— Consider a general $\Lambda$-partition quantum state expressed as $\hat{\rho}_{\Lambda}=\sum_{\mathcal{K}\in\Lambda}P_{\mathcal{K}}\hat{\rho}_{\mathcal{K}}.$ Due to the convexity of QFI, we have
$F_{Q}(\hat{\rho}_{\Lambda},\hat{A})\leq\sum_{\mathcal{K}}P_{\mathcal{K}}F_{Q}(\hat{\rho}_{\mathcal{K}},\hat{A})\leq 4\sum_{\mathcal{K}}P_{\mathcal{K}}\mathcal{V}(\hat{\rho}_{\mathcal{K}},\hat{A}).$ (2)
This can be rewritten as
$\sum_{\mathcal{K}}P_{\mathcal{K}}\left(F_{Q}(\hat{\rho}_{\mathcal{K}},\hat{A})-4\mathcal{V}(\hat{\rho}_{\mathcal{K}},\hat{A})\right)\leq 0.$ (3)
Thus, if there exists an operator $\hat{A}$ for a quantum state $\hat{\rho}$ that can violate Eq. (1) across all partition for different $\mathcal{K}$, then the inequality (3) is violated for any choice of probability $P_{\mathcal{K}}$, implying $\Lambda$-partite inseparability. $\Box$
The most challenging step for entanglement characterization is thus to find the operators $\hat{A}$ that capture the correlations for all partitions. Previous work in CV systems [33, 34] can only identify specific partition, which hinders the understanding of multipartite entanglement structure in CV systems. To this aim, we propose a systematic approach to characterize entanglement structure in the largest possible detection range by analytically optimizing the operators $\hat{A}$ for arbitrary quantum states. The methodology consists of the following steps:
Step 1.— Finding the convex sets of operators witnessing entanglement for each $\mathcal{K}$-partition.
To witness $\mathcal{K}$-partite entanglement through inequality (1) the choice of suitable local operators $\hat{A}_{j}$ is crucial and dependent on the given state $\hat{\rho}$. These operators can be constructed by analytically optimizing over linear combinations of accessible operators forming the set $\mathbf{S}_{j}=\{S_{j}^{(1)},S_{j}^{(2)},\cdots\}$ [39]. Specifically, we express $\hat{A}_{j}$ as the linear combination
$\hat{A}_{j}=\sum_{m=1}c_{j}^{(m)}\hat{S}_{j}^{(m)}=\mathbf{c}_{j}\cdot\mathbf{S}_{j}\;.$ (4)
A typical choice for such a set includes local position operators $\hat{x}_{j}$ and momentum operators $\hat{p}_{j}$. However, since the states we consider are in general non-Gaussian, simply measuring linear observables is insufficient for characterizing their correlations. We therefore extend the family of accessible operators by incorporating higher-order moments of local quadrature operators, e.g. $\mathbf{S}_{j}=[\hat{x}_{j},\hat{p}_{j},\hat{x}_{j}^{2},\hat{p}_{j}^{2},(\hat{x}_{j}\hat{p}_{j}+\hat{p}_{j}\hat{x}_{j})/2]^{T}$. The full operator $\hat{A}(\mathbf{c})$ is then $\hat{A}(\mathbf{c})=\sum_{j=1}^{N}\hat{A}_{j}=\sum_{j=1}^{N}\mathbf{c}_{j}\cdot\mathbf{S}_{j}$. According to Eq. (1), the quantity $W=F_{Q}[\hat{\rho}_{\mathcal{K}},\hat{A}]-4\mathcal{V}(\hat{\rho}_{\mathcal{K}},\hat{A})$ must be nonpositive when the state is separable. To witness entanglement, we maximize $W$ by varying $\mathbf{c}$ to obtain an optimal operator $\hat{A}(\mathbf{c}^{\text{opt}})$. This optimization problem can be formulated mathematically by expressing $W$ in a Rayleigh quotient form $W(\mathcal{M},\mathbf{c})=\frac{\mathbf{c}^{T}\mathcal{M}\mathbf{c}}{\mathbf{c}^{T}\mathbf{c}}$ [33]. Here, $\mathcal{M}$ is a Hermitian matrix associated with the operator set $\mathbf{S}_{j}$ and the targeted quantum states (see the supplementary material [41] for details).
To determine the convex sets of operators for $\mathcal{K}$-partition inseparability, we decompose the matrix $\mathcal{M}$ as follows: $\mathcal{M}=UDU^{T},$ where $U$ is an orthogonal matrix, satisfying $U^{T}U=UU^{T}=I$, and $D$ is a diagonal matrix containing the eigenvalues $\lambda_{i}$ of $\mathcal{M}$. A normalized vector $\mathbf{c}$ can be expressed by the eigenvectors $\mathbf{e}_{i}$: $\mathbf{c}=\sum_{i}n_{i}\mathbf{e}_{i}.$ Substituting this into the Rayleigh quotient $W$, we obtain:
$W(\mathcal{M},\mathbf{c})=$ $\sum_{ij}n_{i}n_{j}\mathbf{e}_{i}^{T}UDU^{T}\mathbf{e}_{j}$ (5)
$=$ $\sum_{ij}n_{i}n_{j}D_{ij}=\sum_{i}n_{i}^{2}\lambda_{i}.$
This indicates that if there is a subset of eigenvectors $Q$, where $\forall\mathbf{e}_{i}\in Q$, $W(\mathcal{M},\mathbf{e}_{i})>0$, then any linear combination $\mathbf{c}=\sum_{i}n_{i}\mathbf{e}_{i}$ will also satisfy $W(\mathcal{M},\mathbf{c})>0$. Therefore, all vectors $\mathbf{e}_{i}$ for which $W(\mathcal{M},\mathbf{e}_{i})>0$ form convex sets. This completes Step 1.
Step 2.—Determining the intersection of convex sets for different types of entanglement structures.
For each $\mathcal{K}$-partition of the $N$ parties, the convex sets of operators determined in Step 1 corresponds to a subspace. Different $\mathcal{K}$-partitions will in general generate different subspaces $Q_{j}$, and the overlap of these subspaces $P=\cap_{j=1}^{m}Q_{j}$ (if any) indicates a common set of operators that allows us to certify $\Lambda$-partition inseparability.
Step 3.—Find the suitable operator to witness entanglement within the intersection space.
The operators for detecting entanglement structure can be optimized by maximizing the minimum value of $W(\mathcal{M}_{i},\mathbf{c})$ across all matrices $\mathcal{M}_{i}$,
$g=\min\{W(\mathcal{M}_{1},\mathbf{c}),W(\mathcal{M}_{2},\mathbf{c}),W(\mathcal{M}_{3},\mathbf{c}),\ldots\},$ (6)
with $\mathcal{M}_{i}$ are associated to each $\mathcal{K}$-partition. However, this optimization can be analytically challenging with the existence of different Rayleigh quotients $W(\mathcal{M}_{i},\mathbf{c})$. To simplify the optimization, we construct a combined function that integrates different $\mathcal{M}_{i}$
$f=$ $\sum_{i}W(\mathcal{M}_{i},\mathbf{c})=W\left(\sum_{i}\mathcal{M}_{i},\mathbf{c}\right)$ (7)
$=$ $\sum_{i=1}^{k}\sum_{j=1}^{k}n_{i}n_{j}(\mathbf{p}_{i}^{T}M\mathbf{p}_{j})=\mathbf{n}^{T}\mathcal{Q}\mathbf{n}\;,$
where we have now the vector $\mathbf{c}=\sum_{i}n_{i}\mathbf{p}_{i}$, with $\mathbf{p}_{i}$ being the basis in the intersection space $P$, and $n_{i}$ are the coefficients to be determined [42]. Here, $\mathcal{Q}$ is a symmetric matrix with elements $q_{ij}=\mathbf{p}_{i}^{T}\mathcal{M}\mathbf{p}_{j}$.
The relatively optimal $\mathbf{c}$ can be found by maximizing the value of $g$ in Eq. (6). This involves selecting the positive eigenvalues $\lambda_{j}$ of $\mathcal{Q}$ that satisfy $g>0$ and then determining the eigenvector corresponding to the optimal $\lambda_{j}$ that maximizes $g$. This ends Step 3.
These three steps provide a general method to witness $\Lambda$-partition inseparability. In the following, we will apply this methodology to various general CV states (both Gaussian and non-Gaussian), demonstrating its effectiveness in characterizing multipartite entanglement structures.
Figure 2: Multipartite entanglement structure detection. Without loss of generality, we randomly generated $10^{4}$ four-mode entangled states for each partition based on Young diagrams. The vertical axis shows the successful rate of detecting this type of entanglement. The color denotes the order of the encoding operators in QFI. The error bars represent the standard deviation of the detection rate, calculated by dividing these $10^{4}$ states into 10 sets.
Characterizing entanglement structure of random Gaussian and non-Gaussian states.— The generation of a random CV quantum state $\hat{\rho}$ begins with a core state $|C\rangle$ using the stellar formalism [43, 44, 45]. A multimode quantum state is then generated by applying an $m$-mode random Gaussian unitary operation $\hat{G}$ to the core state, namely as $|\psi\rangle=\hat{G}|C\rangle$. This procedure results in states that includes the most common Gaussian and non-Gaussian states prepared in experiments, such as squeezed and photon subtracted states (see details in the supplementary material [41]).
As an illustration of the applicability of our entanglement characterization method, we randomly generate $10^{4}$ four-mode quantum states. The entanglement structure of these states cannot be characterized by the previous criteria based on covariance matrix [46, 47, 48, 49, 50] and QFI [33, 34]. The primary limitations of these methods are their inability to detect entanglement across different partitions due to their focus on a specific partition, or their reliance on measuring variances of linear combinations of position and momentum operators, which do not capture non-Gaussian correlations. In contrast, our method based on QFI overcomes these challenges, demonstrating a high success rate in detecting entanglement structures, as shown in Fig. 2. For four modes, using first-order encoding operators $\hat{A}$ (combinations of $\hat{x}$ and $\hat{p}$) we can reach over 90% detection rate. By including higher-order operators ($\hat{x}^{2},\hat{p}^{2},\hat{x}\hat{p}+\hat{p}\hat{x},\dots$), the successful rate of detecting GME increases to almost 100%.
Figure 3: Witnessing GME for non-Gaussian states. (a) The success rate of detecting GME for different multimode cases, where we randomly generated $10^{4}$ genuine entangled states of 3, 4, and 5 modes. The color denotes the order of the encoding operators in QFI. The error bars represent the standard deviation of the percentage of GME, calculated by dividing these $10^{4}$ states into 10 sets. (b) The percentage of detectable GME in the presence of channel loss, where the blue and green denote the QFI criterion based on first-order and second-order operators, respectively. (c) The witness of GME using QFI based on the three-photon SPDC process, where we generated $10^{4}$ states with random nonlinear strengths and plotted the percentage of detectable GME using operators of different orders.
Detecting GME of non-Gaussian states. For $N$ parties to be considered genuinely multipartite entangled, all possible biseparable mixtures must be excluded. Detecting GME of non-Gaussian CV states poses significant challenges due to ruling out all biseparable mixtures. Our method based on QFI can overcome these challenges and can effectively detect GME for non-Gaussian states. As shown in Fig. 3(a), our approach exhibits a high success rate in detecting GME. For three modes, using first-order encoding operators $\hat{A}$ (combinations of $\hat{x}$ and $\hat{p}$), we successfully detect GME in around 94.3% of three-mode cases. By incorporating second-order encoding operators ($\hat{x}^{2},\hat{p}^{2},\hat{x}\hat{p}+\hat{p}\hat{x}$), the successful rate of detecting GME increases to 99.3%. This performance further improves by extending the operator family to higher orders, achieving 99.8% successful rate with fourth-order operators. Additionally, the effectiveness of our method is not confined to three-mode systems; Similar results can also be extended to four and five modes, as illustrated in Fig. 3(a), demonstrating the method’s versatility across different multimode systems.
Furthermore, we evaluate the performance of our method in the presence of a loss channel. Losses are introduced to the pure GME state $|\psi\rangle=\hat{G}|C\rangle$ using a single-mode loss channel $\hat{L}_{i}(\eta)$ with efficiency coefficient $\eta$, as detailed in Ref. [51]. The state with loss is described by: $\hat{\rho}=\left(\prod_{i=1}^{m}\hat{L}_{i}(\eta_{i})\right)\hat{G}|C\rangle\langle C|\hat{G}^{\dagger}\left(\prod_{i=1}^{m}\hat{L}_{i}^{\dagger}(\eta_{i})\right).$ Increased channel loss ($1-\eta$) impacts both the entanglement itself and the detection of GME. As indicated by the blue curve in Fig. 3(b), the percentage of GME detectable using first-order QFI decreases with increasing loss. Intriguingly, we demonstrate that extending the analysis to include second-order operators significantly enhances the successful detection rate of GME, showing robustness against channel loss. This improvement is depicted in the green curves of Fig. 3(b).
In addition to examining non-Gaussian states generated from core states, we also explore non-Gaussian states produced by nonlinear processes, such as the three-photon SPDC process. This process has been realized experimentally [27, 26] and has theoretically attracted significant interest due to its challenges for characterizing its GME [52, 53, 54, 55, 56, 57]. For our analysis, we randomly generated $10^{4}$ non-Gaussian states using the three-photon SPDC process, described by the Hamiltonian $\hat{H}=(\chi_{1}\hat{a}\hat{b}^{2}+\chi_{2}\hat{b}\hat{c}^{2}+\chi_{3}\hat{c}\hat{a}^{2})+h.c.$ Here the parameters $\chi_{i}t$ are randomly selected from the interval [0, 0.04], and the initial states are vacuum states. As illustrated in Fig. 3(c), the specific form of the three-photon Hamiltonian means that the first-order QFI fails to detect any quantum correlations. In contrast, higher-order QFI successfully identifies GME. These results exhibit the versatility of our method across different non-Gaussian states, demonstrating its effectiveness in accurately identifying operators that capture complex non-Gaussian correlations.
In conclusion, we provide a general method for characterizing entanglement for arbitrarily Gaussian and non-Gaussian states. Based on QFI, we establish a criterion and propose a systematic approach to identify entanglement by analytically optimizing operators from accessible family of operators. This method involves three key steps: (i) identifying convex sets of operators that can be used to witness $\mathcal{K}$-partition entanglement, (ii) determining the intersection of these sets, and (iii) selecting a suitable operator at the intersection. To verify the accessibility and effectiveness of our method, we randomly generated $10^{4}$ states based on core states and nonlinear processes, which demonstrate a very high proportion of success in detecting entanglement structure. Furthermore, we exhibit the advantage of our method in resisting channel loss and detecting non-Gaussian GME. This approach, for the first time, provides a systematic way to distinguish entanglement structure for arbitrary states in CV systems, paving the way for further exploration and detection of entanglement in multimode experimental systems.
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (No. 12125402), the Innovation Program for Quantum Science and Technology (No. 2021ZD0301500), and Beijing Natural Science Foundation (Grant No. Z240007). M.F. was supported by the Swiss National Science Foundation Ambizione Grant No. 208886, and by The Branco Weiss Fellowship – Society in Science, administered by the ETH Zürich.

### v2

@@
-Characterization of Multipartite non-Gaussian Entanglement Structure
+Metrological Characterization of Multipartite Continuous-Variable non-Gaussian Entanglement Structure
 Abstract
 Multipartite entanglement is an essential resource for quantum information tasks, but characterizing entanglement structures in continuous variable systems remains challenging, especially in multimode non-Gaussian scenarios.
-In this work, we introduce a method for detecting multipartite entanglement structures in arbitrary continuous variable states, which also reveals genuine multipartite entanglement.
-By leveraging the quantum Fisher information, we propose a criterion for characterizing entanglement and offer a systematic approach to identify feasible operators that capture quantum correlations in multimode non-Gaussian states.
-We demonstrate the effectiveness of our method on over $10^{4}$ randomly generated multimode-entangled quantum states, achieving a high success rate in entanglement detection.
+In this work, we introduce a method for detecting multipartite entanglement structures in continuous variable states.
+By leveraging the quantum Fisher information, we propose a systematic approach to identify feasible operators that capture quantum correlations in multimode non-Gaussian states.
+We demonstrate the effectiveness of our method on over $10^{5}$ randomly generated multimode-entangled quantum states, achieving a high success rate in entanglement detection.
 Additionally, our method exhibits enhanced robustness against losses by expanding the set of accessible operators.
@@
 Its importance is particularly evident in continuous variable (CV) systems, where multimode Gaussian states, such as squeezed states and cluster states, can be deterministically prepared and controlled [11, 12, 13].
-Moreover, non-Gaussian features in CV systems are proven to be a resource providing an advantage over classical protocols [14].
-These non-Gaussian states play a crucial role for quantum computation [15, 16, 17], sensing [18, 19, 20, 21], and imaging [22, 23].
-Strategies to generate and control such states has seen substantial advancements in recent years, including through photon-subtraction [24, 25], spontaneous parametric down-conversion (SPDC) [26, 27], as well as in microwave cavities coupled to superconducting artificial atoms [28, 29].
-Despite these advancements, accurately characterizing both Gaussian and non-Gaussian multimode entanglement in CV systems continues to present substantial challenges [14], e.g.
+However, to achieve an advantage over classical protocols in CV systems, non-Gaussian features are necessary [14].
+Non-Gaussian states play a crucial role for quantum computation [15, 16, 17], sensing [18, 19, 20, 21], and imaging [22, 23].
+Strategies to generate and control such states has seen substantial advancements in recent years, including through photon-subtraction [24, 25] and nonlinear operations, such as spontaneous parametric down-conversion (SPDC) [26, 27] and Kerr interactions in microwave cavities coupled superconducting artificial atoms [28, 29, 30].
+Despite these advancements, accurately characterizing both Gaussian and non-Gaussian multimode entanglement in CV systems continues to present significant challenges [14], e.g.
 due to the complex correlations appearing among higher-order statistical moments of quadrature field operations.
-Unlike discrete variable (DV) systems, where entanglement structures are identified based on separable partitions and their sizes [30, 31, 32], there is no such comprehensive method for CV systems.
-Conventional works [33, 34, 35] in CV systems can only distinguish entanglement structures for the one given partition and become ineffective with convex combinations of different partitions.
-This limitation hinders the understanding of complex structures, such as m-mode genuine multipartite entanglement.
-Moreover, while quantum correlations in Gaussian states are predominantly analyzed through first and second-order moments [36, 34, 37], non-Gaussian states involve more complex higher-order correlations.
-In general, these correlations can not be captured by standard criteria based on Gaussian metrics, which typically focus on the first- and second-order correlations alone.
-Therefore, there is a pressing need to develop a method to characterize entanglement structure in CV systems, filling this significant gap in the field.
+Unlike discrete variable (DV) systems, where entanglement structures are identified based on separable partitions and their sizes [31, 32, 33, 34, 35], there is no similarly comprehensive method for CV systems.
+State-of-the-art techniques in CV systems predominantly analyze bi- and multipartite entanglement through first and second-order moments [36, 37, 38, 39, 40, 41, 42, 43, 44].
+Since non-Gaussian states involve more complex higher-order correlations, these techniques are mostly limited to the study of Gaussian CV entanglement.
+Criteria for detecting non-Gaussian CV entanglement [45, 46, 47, 48, 49] can only distinguish entanglement properties of individual partitions and become ineffective with different partitions.
+This limitation hinders the understanding of complex structures, such as the fully inseparable entanglement structure.
+Therefore, there is a pressing need to develop a method to characterize the multipartite entanglement structure in non-Gaussian CV systems, filling this significant gap in the field.
 Here, we propose a general method to characterize entanglement structure based on quantum metrology tools.
@@
 Since different entanglement structures correlate with distinct convex sets, we then proceed to pinpoint the intersection of these sets.
-This intersection allows us to identify a relatively optimal operator for witnessing entanglement.
-Crucially, contrary to previous work which allows for identifying a specific entanglement structure, our method allows for identifying also their convex combinations, which enables the classification of entanglement and detection of genuine multipartite entanglement (GME).
-To demonstrate the efficiency of our method, we apply it to a broad range of non-Gaussian states, generated through various stellar ranks and nonlinear processes, including over $10^{4}$ randomly produced quantum states.
-Additionally, we examine the robustness of our method in scenarios involving loss, demonstrating that the inclusion of high-order operators significantly enhances resilience to loss channels.
-This approach marks the first accessible method for witnessing entanglement structure in arbitrary Gaussian and non-Gaussian states, thereby addressing a significant need in CV systems.
+This intersection allows us to identify an optimal operator for witnessing entanglement.
+Crucially, contrary to previous work where the chosen operator can identify a specific entanglement structure, our method identifies different entanglement structures with the optimal operator, which enables us to efficiently characterize and detect entanglement structures.
+To demonstrate the practicality and efficiency of our method, we apply it to a broad range of non-Gaussian states generated through various stellar ranks and nonlinear processes, including over $10^{5}$ randomly produced quantum states.
+Furthermore, we examine the robustness of our method in scenarios involving loss, demonstrating that the inclusion of high-order operators significantly enhances resilience to loss channels.
+This approach marks the accessible method for witnessing entanglement structure in arbitrary Gaussian and non-Gaussian states, thereby addressing a significant need in CV systems.
 Figure 1: Graph representation of different multipartite entanglement structures.
 As illustrated, a system with $N=4$ modes can be characterized by 15 different partitions, associated to different distributions of correlations as indicated by the solid circles.
-Each partition can be graphically represented by a Young diagram, where the number of boxes in each row indicates the largest of correlated modes, while the number of boxes in the first column indicates the number of partitions of the system.
-Metrological bound for $\Lambda$-partitions.– We begin discussing the entanglement structure for $N$ parties.
+Each partition can be graphically represented by a Young tableau, where boxes in each row indicate correlated modes, while vertically stacked boxes indicate separable subsets.
+Metrological bound for $\mathcal{K}$-partitions.– We begin discussing the entanglement structure for $N$ parties.
 A partition $\mathcal{K}=\left\{H_{1},H_{2},\ldots,H_{k}\right\}$ separates the total $N$-partite system into $k$ nonempty, disjoint subsets $H_{l}$, where each subset has a size $N_{l}$ such that $\sum_{l=1}^{k}N_{l}=N$.
-A state $\hat{\rho}_{\mathcal{K}}$ is $\mathcal{K}$-separable if there exist local quantum states $\hat{\rho}_{H_{l}}^{(\gamma)}$ for each subset and a probability distribution $p_{\gamma}$ such that $\hat{\rho}_{\mathcal{K}}=\sum_{\gamma}p_{\gamma}\hat{\rho}_{H_{1}}^{(\gamma)}\otimes\cdots\otimes\hat{\rho}_{H_{k}}^{(\gamma)}$.
-A partition $\mathcal{K}$ can be classified according to the number of subsets and the (largest) size of the subsets.
-This can be illustrated graphically by means of Young diagrams, as illustrated in Fig.
+A state $\hat{\rho}_{\mathcal{K}}$ is $\mathcal{K}$-separable if there exists local quantum states $\hat{\rho}_{H_{l}}^{(\gamma)}$ for each subset and a probability distribution $p_{\gamma}$ such that $\hat{\rho}_{\mathcal{K}}=\sum_{\gamma}p_{\gamma}\hat{\rho}_{H_{1}}^{(\gamma)}\otimes\cdots\otimes\hat{\rho}_{H_{k}}^{(\gamma)}$.
+A partition $\mathcal{K}$ can be classified in terms of the number and respective sizes of the separable subsets.
+This can be illustrated graphically by means of Young tableau, as illustrated in Fig.
 1 for a 4-mode system.
-We employ the notation $\Lambda$ to refer to convex combinations of partitions associated with identical Young diagrams.
-And any states that cannot be written as a $\Lambda$-separable state $\hat{\rho}_{\Lambda}=\sum_{\mathcal{K}\in\Lambda}P_{\mathcal{K}}\hat{\rho}_{\mathcal{K}}$ are denoted as $\Lambda$-inseparable states [38].
 Using the framework of quantum metrology, it is known that for an arbitrary $\mathcal{K}$-separable state $\hat{\rho}_{\mathcal{K}}$, the quantum Fisher information (QFI) is bounded by the variances as [46, 43]
-$F_{Q}[\hat{\rho}_{\mathcal{K}},\hat{A}]\leq 4\mathcal{V}(\hat{\rho}_{\mathcal{K}},\hat{A}).$ (1)
-Here, $\hat{A}=\sum_{j=1}^{N}\hat{A}_{j}$ with $\hat{A}_{j}$ represents a local generator in the reduced state $\hat{\rho}_{j}$, $\mathcal{V}(\hat{\rho}_{\mathcal{K}},\hat{A})=\sum_{i=1}^{k}\mathrm{Var}(\hat{\rho}_{H_{i}},\sum_{j\in H_{i}}\hat{A}_{j})$, and $\mathrm{Var}(\hat{\rho},\hat{O})=\langle\hat{O}^{2}\rangle_{\hat{\rho}}-\langle\hat{O}\rangle^{2}_{\hat{\rho}}$ defines the variance.
-The explicit expression of QFI is given by $F_{Q}[\hat{\rho}_{\mathcal{K}},\hat{A}]=2\sum_{\begin{subarray}{c}k,l\\ p_{k}+p_{l}\neq 0\end{subarray}}\frac{(p_{k}-p_{l})^{2}}{p_{k}+p_{l}}|\langle\Psi_{k}|\hat{A}|\Psi_{l}\rangle|^{2}$, with the spectral decomposition $\hat{\rho}=\sum_{k}p_{k}|\Psi_{k}\rangle\langle\Psi_{k}|$ [40].
+$F_{Q}(\hat{\rho}_{\mathcal{K}},\hat{A})\leq 4\mathcal{V}(\hat{\rho}_{\mathcal{K}},\hat{A}).$ (1)
+Here, $\hat{A}=\sum_{j=1}^{N}\hat{A}_{j}$ with $\hat{A}_{j}$ represents a local generator in the reduced state $\hat{\rho}_{j}$, $\mathcal{V}(\hat{\rho}_{\mathcal{K}},\hat{A})=\sum_{i=1}^{k}\mathrm{Var}(\hat{\rho}_{H_{i}},\sum_{j\in H_{i}}\hat{A}_{j})$, $\mathrm{Var}(\hat{\rho},\hat{O})=\langle\hat{O}^{2}\rangle_{\hat{\rho}}-\langle\hat{O}\rangle^{2}_{\hat{\rho}}$ defines the variance, and $\hat{\rho}_{H_{i}}=\sum_{\gamma}p_{\gamma}\hat{\rho}_{H_{i}}^{(\gamma)}$ the reduced state of subsystem $H_{i}$.
+The explicit expression of the QFI is given by $F_{Q}[\hat{\rho}_{\mathcal{K}},\hat{A}]=2\sum_{\begin{subarray}{c}k,l\\ p_{k}+p_{l}\neq 0\end{subarray}}\frac{(p_{k}-p_{l})^{2}}{p_{k}+p_{l}}|\langle\Psi_{k}|\hat{A}|\Psi_{l}\rangle|^{2}$, with the spectral decomposition $\hat{\rho}=\sum_{k}p_{k}|\Psi_{k}\rangle\langle\Psi_{k}|$ [50].
 Since Eq.
-(1) is a necessary criterion for separability, its violation implies $\mathcal{K}$-paritite inseparability.
-To distinguish different $\Lambda$ partitions, a general metrological bound that is suitable for all mixed $\mathcal{K}$-separable states with $\mathcal{K}\in\Lambda$ needs to be given.
-Theorem 1. $\Lambda$-partite inseparability can be confirmed if there is at least one operator $\hat{A}$ that can violate Eq. (1) across all $\mathcal{K}$-partition with $\mathcal{K}\in\Lambda$.
-Proof.— Consider a general $\Lambda$-partition quantum state expressed as $\hat{\rho}_{\Lambda}=\sum_{\mathcal{K}\in\Lambda}P_{\mathcal{K}}\hat{\rho}_{\mathcal{K}}.$ Due to the convexity of QFI, we have
-$F_{Q}(\hat{\rho}_{\Lambda},\hat{A})\leq\sum_{\mathcal{K}}P_{\mathcal{K}}F_{Q}(\hat{\rho}_{\mathcal{K}},\hat{A})\leq 4\sum_{\mathcal{K}}P_{\mathcal{K}}\mathcal{V}(\hat{\rho}_{\mathcal{K}},\hat{A}).$ (2)
-This can be rewritten as
-$\sum_{\mathcal{K}}P_{\mathcal{K}}\left(F_{Q}(\hat{\rho}_{\mathcal{K}},\hat{A})-4\mathcal{V}(\hat{\rho}_{\mathcal{K}},\hat{A})\right)\leq 0.$ (3)
-Thus, if there exists an operator $\hat{A}$ for a quantum state $\hat{\rho}$ that can violate Eq.
-(1) across all partition for different $\mathcal{K}$, then the inequality (3) is violated for any choice of probability $P_{\mathcal{K}}$, implying $\Lambda$-partite inseparability.
-$\Box$
-The most challenging step for entanglement characterization is thus to find the operators $\hat{A}$ that capture the correlations for all partitions.
-Previous work in CV systems [33, 34] can only identify specific partition, which hinders the understanding of multipartite entanglement structure in CV systems.
-To this aim, we propose a systematic approach to characterize entanglement structure in the largest possible detection range by analytically optimizing the operators $\hat{A}$ for arbitrary quantum states.
+(1) is a necessary criterion for separability, its violation implies $\mathcal{K}$-inseparability.
+The right-hand side of the criterion (1) is tailored to detect inseparability within the specific partition $\mathcal{K}$.
+To maximize entanglement detection, one must optimize the operator $A$ to achieve the largest possible violation of Eq.
+(1).
+Different partitions typically correspond to different operators $\hat{A}$.
+Therefore, if an operator $\hat{A}$ exists for a quantum state $\hat{\rho}$ that can capture correlations across all partitions corresponding to a given Young diagram, as illustrated in Fig.
+1, it offers a more efficient way to characterize the entanglement structure.
+To this aim, we propose a general and systematic optimization method to characterize different partitions with the same operator $\hat{A}$.
 The methodology consists of the following steps:
@@
 Here, $\mathcal{Q}$ is a symmetric matrix with elements $q_{ij}=\mathbf{p}_{i}^{T}\mathcal{M}\mathbf{p}_{j}$.
-The relatively optimal $\mathbf{c}$ can be found by maximizing the value of $g$ in Eq.
+The optimal $\mathbf{c}$ can be found by maximizing the value of $g$ in Eq.
 (4).
@@
 This ends Step 3.
-These three steps provide a general method to witness $\Lambda$-partition inseparability.
+These three steps provide a general method for finding an operator to witness entanglement for different partitions.
 In the following, we will apply this methodology to various general CV states (both Gaussian and non-Gaussian), demonstrating its effectiveness in characterizing multipartite entanglement structures.
 Figure 2: Multipartite entanglement structure detection.
-Without loss of generality, we randomly generated $10^{4}$ four-mode entangled states for each partition based on Young diagrams.
+Without loss of generality, we randomly generated $10^{4}$ four-mode entangled states for each entanglement structure, classified according to Young diagram.
 The vertical axis shows the successful rate of detecting this type of entanglement.
-The color denotes the order of the encoding operators in QFI.
+The color denotes the order of the encoding operators $\mathbf{S}_{j}$ in QFI.
 The error bars represent the standard deviation of the detection rate, calculated by dividing these $10^{4}$ states into 10 sets.
@@
 This procedure results in states that includes the most common Gaussian and non-Gaussian states prepared in experiments, such as squeezed and photon subtracted states (see details in the supplementary material [51]).
-As an illustration of the applicability of our entanglement characterization method, we randomly generate $10^{4}$ four-mode quantum states.
-The entanglement structure of these states cannot be characterized by the previous criteria based on covariance matrix [46, 47, 48, 49, 50] and QFI [33, 34].
+The entanglement structure of these random generated multimode states cannot be characterized by the previous criteria based on covariance matrix [38, 39, 40, 36, 37] and QFI [43, 47].
 The primary limitations of these methods are their inability to detect entanglement across different partitions due to their focus on a specific partition, or their reliance on measuring variances of linear combinations of position and momentum operators, which do not capture non-Gaussian correlations.
-In contrast, our method based on QFI overcomes these challenges, demonstrating a high success rate in detecting entanglement structures, as shown in Fig.
-2.
+In contrast, our method based on QFI overcomes these challenges.
+To test the effectiveness of our method, We randomly generate $40000$ four-mode quantum states (ensuring 10,000 states for each type of entanglement structure).
+As shown in Fig.
+2, our method based on QFI demonstrating a high success rate in detecting entanglement structures.
 For four modes, using first-order encoding operators $\hat{A}$ (combinations of $\hat{x}$ and $\hat{p}$) we can reach over 90% detection rate.
-By including higher-order operators ($\hat{x}^{2},\hat{p}^{2},\hat{x}\hat{p}+\hat{p}\hat{x},\dots$), the successful rate of detecting GME increases to almost 100%.
-Figure 3: Witnessing GME for non-Gaussian states.
-(a) The success rate of detecting GME for different multimode cases, where we randomly generated $10^{4}$ genuine entangled states of 3, 4, and 5 modes.
+By including higher-order operators ($\hat{x}^{2},\hat{p}^{2},\hat{x}\hat{p}+\hat{p}\hat{x},\dots$), the successful rate of characterizing entanglement structure increases to almost 100%.
+Figure 3: Witnessing fully inseparable entanglement for different non-Gaussian states.
+(a) The success rate of detecting entanglement for different multimode cases, where we randomly generated $10^{4}$ fully inseparable states of 3, 4, and 5 modes.
 The color denotes the order of the encoding operators in QFI.
-The error bars represent the standard deviation of the percentage of GME, calculated by dividing these $10^{4}$ states into 10 sets.
-(b) The percentage of detectable GME in the presence of channel loss, where the blue and green denote the QFI criterion based on first-order and second-order operators, respectively.
-(c) The witness of GME using QFI based on the three-photon SPDC process, where we generated $10^{4}$ states with random nonlinear strengths and plotted the percentage of detectable GME using operators of different orders.
-Detecting GME of non-Gaussian states.
-For $N$ parties to be considered genuinely multipartite entangled, all possible biseparable mixtures must be excluded.
-Detecting GME of non-Gaussian CV states poses significant challenges due to ruling out all biseparable mixtures.
-Our method based on QFI can overcome these challenges and can effectively detect GME for non-Gaussian states.
-As shown in Fig.
-3(a), our approach exhibits a high success rate in detecting GME.
-For three modes, using first-order encoding operators $\hat{A}$ (combinations of $\hat{x}$ and $\hat{p}$), we successfully detect GME in around 94.3% of three-mode cases.
-By incorporating second-order encoding operators ($\hat{x}^{2},\hat{p}^{2},\hat{x}\hat{p}+\hat{p}\hat{x}$), the successful rate of detecting GME increases to 99.3%.
+The error bars represent the standard deviation of the percentage of detecting fully inseparable states, calculated by dividing these $10^{4}$ states into 10 sets.
+(b) The percentage of detectable fully inseparable states in the presence of channel loss, where the blue and green denote the QFI criterion based on first-order and second-order operators, respectively.
+(c) The witness of full inseparability using QFI based on the three-photon SPDC process, where we generated $10^{4}$ states with random nonlinear strengths and plotted the percentage of detectable full inseparable states using operators of different orders.
+Additionally, the effectiveness of our method is not confined to four-mode systems; Similar results can also be extended to four and five modes.
+As illustrated in Fig.
+3(a), our approach exhibits a high success rate in detecting fully inseparable in different multimodes.
+For three modes, using first-order encoding operators $\hat{A}$ (combinations of $\hat{x}$ and $\hat{p}$), we successfully detect fully inseparable entanglement in around 94.3% of three-mode cases.
+By incorporating second-order encoding operators ($\hat{x}^{2},\hat{p}^{2},\hat{x}\hat{p}+\hat{p}\hat{x}$), the successful detecting rate increases to 99.3%.
 This performance further improves by extending the operator family to higher orders, achieving 99.8% successful rate with fourth-order operators.
-Additionally, the effectiveness of our method is not confined to three-mode systems; Similar results can also be extended to four and five modes, as illustrated in Fig.
-3(a), demonstrating the method’s versatility across different multimode systems.
 Furthermore, we evaluate the performance of our method in the presence of a loss channel.
-Losses are introduced to the pure GME state $|\psi\rangle=\hat{G}|C\rangle$ using a single-mode loss channel $\hat{L}_{i}(\eta)$ with efficiency coefficient $\eta$, as detailed in Ref.
+Losses are introduced to the pure entangled state $|\psi\rangle=\hat{G}|C\rangle$ using a single-mode loss channel $\hat{L}_{i}(\eta)$ with efficiency coefficient $\eta$, as detailed in Ref.
 [56].
-The state with loss is described by: $\hat{\rho}=\left(\prod_{i=1}^{m}\hat{L}_{i}(\eta_{i})\right)\hat{G}|C\rangle\langle C|\hat{G}^{\dagger}\left(\prod_{i=1}^{m}\hat{L}_{i}^{\dagger}(\eta_{i})\right).$ Increased channel loss ($1-\eta$) impacts both the entanglement itself and the detection of GME.
+The state with loss is described by: $\hat{\rho}=\left(\prod_{i=1}^{m}\hat{L}_{i}(\eta_{i})\right)\hat{G}|C\rangle\langle C|\hat{G}^{\dagger}\left(\prod_{i=1}^{m}\hat{L}_{i}^{\dagger}(\eta_{i})\right).$ Increased channel loss ($1-\eta$) impacts both the entanglement itself and the successful detection rate.
 As indicated by the blue curve in Fig.
-3(b), the percentage of GME detectable using first-order QFI decreases with increasing loss.
-Intriguingly, we demonstrate that extending the analysis to include second-order operators significantly enhances the successful detection rate of GME, showing robustness against channel loss.
+3(b), the percentage of detectable fully inseparable states using first-order QFI decreases with increasing loss.
+Intriguingly, we demonstrate that extending the analysis to include second-order operators significantly enhances the successful detection rate, showing robustness against channel loss.
 This improvement is depicted in the green curves of Fig.
@@
 In addition to examining non-Gaussian states generated from core states, we also explore non-Gaussian states produced by nonlinear processes, such as the three-photon SPDC process.
-This process has been realized experimentally [27, 26] and has theoretically attracted significant interest due to its challenges for characterizing its GME [52, 53, 54, 55, 56, 57].
+This process has been realized experimentally [27, 26] and has theoretically attracted significant interest due to its challenges in characterizing its correlations [57, 58, 59, 60, 61, 62].
 For our analysis, we randomly generated $10^{4}$ non-Gaussian states using the three-photon SPDC process, described by the Hamiltonian $\hat{H}=(\chi_{1}\hat{a}\hat{b}^{2}+\chi_{2}\hat{b}\hat{c}^{2}+\chi_{3}\hat{c}\hat{a}^{2})+h.c.$ Here the parameters $\chi_{i}t$ are randomly selected from the interval [0, 0.04], and the initial states are vacuum states.
@@
 3(c), the specific form of the three-photon Hamiltonian means that the first-order QFI fails to detect any quantum correlations.
-In contrast, higher-order QFI successfully identifies GME.
+In contrast, higher-order QFI successfully identifies entanglement.
 These results exhibit the versatility of our method across different non-Gaussian states, demonstrating its effectiveness in accurately identifying operators that capture complex non-Gaussian correlations.
@@
 This method involves three key steps: (i) identifying convex sets of operators that can be used to witness $\mathcal{K}$-partition entanglement, (ii) determining the intersection of these sets, and (iii) selecting a suitable operator at the intersection.
-To verify the accessibility and effectiveness of our method, we randomly generated $10^{4}$ states based on core states and nonlinear processes, which demonstrate a very high proportion of success in detecting entanglement structure.
-Furthermore, we exhibit the advantage of our method in resisting channel loss and detecting non-Gaussian GME.
-This approach, for the first time, provides a systematic way to distinguish entanglement structure for arbitrary states in CV systems, paving the way for further exploration and detection of entanglement in multimode experimental systems.
+To verify the accessibility and effectiveness of our method, we randomly generated $10^{5}$ states based on core states and nonlinear processes, which demonstrate a very high proportion of success in detecting entanglement structure.
+Furthermore, we exhibit the advantage of our method in resisting channel loss and detecting non-Gaussian entanglement.
+This approach provides a systematic way to distinguish entanglement structure for arbitrary states in CV systems, paving the way for further exploration and detection of entanglement in multimode experimental systems.
 Acknowledgements.
 This work is supported by the National Natural Science Foundation of China (No.
-12125402), the Innovation Program for Quantum Science and Technology (No.
+12125402 and No.
+12474256), the Innovation Program for Quantum Science and Technology (No.
 2021ZD0301500), and Beijing Natural Science Foundation (Grant No.

### v3

@@
-Metrological Characterization of Multipartite Continuous-Variable non-Gaussian Entanglement Structure
+Characterizing the Multipartite Entanglement Structure of Non-Gaussian Continuous-Variable States with a Single Evolution Operator
 Abstract
 Multipartite entanglement is an essential resource for quantum information tasks, but characterizing entanglement structures in continuous variable systems remains challenging, especially in multimode non-Gaussian scenarios.
-In this work, we introduce a method for detecting multipartite entanglement structures in continuous variable states.
-By leveraging the quantum Fisher information, we propose a systematic approach to identify feasible operators that capture quantum correlations in multimode non-Gaussian states.
-We demonstrate the effectiveness of our method on over $10^{5}$ randomly generated multimode-entangled quantum states, achieving a high success rate in entanglement detection.
-Additionally, our method exhibits enhanced robustness against losses by expanding the set of accessible operators.
+In this work, we introduce an efficient method for detecting multipartite entanglement structures in continuous-variable states.
+Based on the quantum Fisher information, we propose a systematic approach to identify an optimal encoding operator that can capture the quantum correlations in multimode non-Gaussian states.
+We demonstrate the effectiveness of our method on over $10^{5}$ randomly generated multimode-entangled quantum states, achieving a very high success rate in entanglement detection.
+Additionally, the robustness of our method can be considerably enhanced against losses by expanding the set of accessible operators.
 This work provides a general framework for characterizing entanglement structures in diverse continuous variable systems, enabling a number of experimentally relevant applications.
@@
 However, to achieve an advantage over classical protocols in CV systems, non-Gaussian features are necessary [14].
-Non-Gaussian states play a crucial role for quantum computation [15, 16, 17], sensing [18, 19, 20, 21], and imaging [22, 23].
+Non-Gaussian states play a crucial role for quantum computation [15, 16, 17], sensing [18, 19, 20], and imaging [21, 22].
 Strategies to generate and control such states has seen substantial advancements in recent years, including through photon-subtraction [23, 24] and nonlinear operations, such as spontaneous parametric down-conversion (SPDC) [25, 26] and Kerr interactions in microwave cavities coupled superconducting artificial atoms [27, 28, 29].
@@
 Unlike discrete variable (DV) systems, where entanglement structures are identified based on separable partitions and their sizes [30, 31, 32, 33, 34], there is no similarly comprehensive method for CV systems.
-State-of-the-art techniques in CV systems predominantly analyze bi- and multipartite entanglement through first and second-order moments [36, 37, 38, 39, 40, 41, 42, 43, 44].
+State-of-the-art techniques in CV systems predominantly analyze bi- and multipartite entanglement through first and second-order moments [35, 36, 37, 38, 39, 40, 41, 42].
 Since non-Gaussian states involve more complex higher-order correlations, these techniques are mostly limited to the study of Gaussian CV entanglement.
-Criteria for detecting non-Gaussian CV entanglement [45, 46, 47, 48, 49] can only distinguish entanglement properties of individual partitions and become ineffective with different partitions.
-This limitation hinders the understanding of complex structures, such as the fully inseparable entanglement structure.
-Therefore, there is a pressing need to develop a method to characterize the multipartite entanglement structure in non-Gaussian CV systems, filling this significant gap in the field.
+Criteria for detecting non-Gaussian CV entanglement [43, 44, 45, 46, 47, 48, 49] can only distinguish entanglement properties of individual partitions and need to use different operators to capture the correlations for other partitions, which become experimentally ineffective to characterize entanglement structure.
+This limitation also hinders the understanding of complex structures.
+Therefore, there is a pressing need to develop an experimentally friendly method to characterize the multipartite entanglement structure in non-Gaussian CV systems, filling the gap in the field.
 Here, we propose a general method to characterize entanglement structure based on quantum metrology tools.
-Specifically, we systematically determine the operators that can effectively capture complex correlations in multimode systems.
+Specifically, we systematically determine an optimal operator that can effectively capture different correlations in multimode systems.
 Our approach begins with the analytical identification of a convex set of operators, each capable of witnessing specific bi-inseparable entanglement structures across arbitrary quantum states.
@@
 This intersection allows us to identify an optimal operator for witnessing entanglement.
-Crucially, contrary to previous work where the chosen operator can identify a specific entanglement structure, our method identifies different entanglement structures with the optimal operator, which enables us to efficiently characterize and detect entanglement structures.
+Crucially, contrary to previous work where one needs to choose different operators to identify different quantum correlations, our method identifies various quantum correlations with the same operator, which enables us to efficiently characterize and detect entanglement structures.
 To demonstrate the practicality and efficiency of our method, we apply it to a broad range of non-Gaussian states generated through various stellar ranks and nonlinear processes, including over $10^{5}$ randomly produced quantum states.
@@
 A partition $\mathcal{K}$ can be classified in terms of the number and respective sizes of the separable subsets.
-This can be illustrated graphically by means of Young tableau, as illustrated in Fig.
+This can be illustrated graphically by means of a Young tableau, as illustrated in Fig.
 1 for a 4-mode system.
-Using the framework of quantum metrology, it is known that for an arbitrary $\mathcal{K}$-separable state $\hat{\rho}_{\mathcal{K}}$, the quantum Fisher information (QFI) is bounded by the variances as [46, 43]
+Using the framework of quantum metrology, it is known that for an arbitrary $\mathcal{K}$-separable state $\hat{\rho}_{\mathcal{K}}$, the quantum Fisher information (QFI) is bounded by the variances as [44]
 $F_{Q}(\hat{\rho}_{\mathcal{K}},\hat{A})\leq 4\mathcal{V}(\hat{\rho}_{\mathcal{K}},\hat{A}).$ (1)
@@
 The right-hand side of the criterion (1) is tailored to detect inseparability within the specific partition $\mathcal{K}$.
-To maximize entanglement detection, one must optimize the operator $A$ to achieve the largest possible violation of Eq.
+To maximize entanglement detection, one must optimize the operator $\hat{A}$ to achieve the largest possible violation of Eq.
 (1).
-Different partitions typically correspond to different operators $\hat{A}$.
+Typically, a different operator must be chosen in order to efficiently witness the entanglement across a different partition.
 Therefore, if an operator $\hat{A}$ exists for a quantum state $\hat{\rho}$ that can capture correlations across all partitions corresponding to a given Young diagram, as illustrated in Fig.
@@
 To witness $\mathcal{K}$-partite entanglement through inequality (1) the choice of suitable local operators $\hat{A}_{j}$ is crucial and dependent on the given state $\hat{\rho}$.
-These operators can be constructed by analytically optimizing over linear combinations of accessible operators forming the set $\mathbf{S}_{j}=\{S_{j}^{(1)},S_{j}^{(2)},\cdots\}$ [46].
+These operators can be constructed by analytically optimizing over linear combinations of accessible operators forming the set $\mathbf{S}_{j}=\{\hat{S}_{j}^{(1)},\hat{S}_{j}^{(2)},\cdots\}$ [44].
 Specifically, we express $\hat{A}_{j}$ as the linear combination
@@
 We therefore extend the family of accessible operators by incorporating higher-order moments of local quadrature operators, e.g.
-$\mathbf{S}_{j}=[\hat{x}_{j},\hat{p}_{j},\hat{x}_{j}^{2},\hat{p}_{j}^{2},(\hat{x}_{j}\hat{p}_{j}+\hat{p}_{j}\hat{x}_{j})/2]^{T}$.
+$\mathbf{S}_{j}=\left(\hat{x}_{j},\hat{p}_{j},\hat{x}_{j}^{2},\hat{p}_{j}^{2},(\hat{x}_{j}\hat{p}_{j}+\hat{p}_{j}\hat{x}_{j})/2\right)$.
 The full operator $\hat{A}(\mathbf{c})$ is then $\hat{A}(\mathbf{c})=\sum_{j=1}^{N}\hat{A}_{j}=\sum_{j=1}^{N}\mathbf{c}_{j}\cdot\mathbf{S}_{j}$.
@@
 To witness entanglement, we maximize $W$ by varying $\mathbf{c}$ to obtain an optimal operator $\hat{A}(\mathbf{c}^{\text{opt}})$.
-This optimization problem can be formulated mathematically by expressing $W$ in a Rayleigh quotient form $W(\mathcal{M},\mathbf{c})=\frac{\mathbf{c}^{T}\mathcal{M}\mathbf{c}}{\mathbf{c}^{T}\mathbf{c}}$ [43].
+This optimization problem can be formulated mathematically by expressing $W$ in a Rayleigh quotient form $W(\mathcal{M},\mathbf{c})=\frac{\mathbf{c}^{T}\mathcal{M}\mathbf{c}}{\mathbf{c}^{T}\mathbf{c}}$.
 Here, $\mathcal{M}$ is a Hermitian matrix associated with the operator set $\mathbf{S}_{j}$ and the targeted quantum states (see the supplementary material [51] for details).
 To determine the convex sets of operators for $\mathcal{K}$-partition inseparability, we decompose the matrix $\mathcal{M}$ as follows: $\mathcal{M}=UDU^{T},$ where $U$ is an orthogonal matrix, satisfying $U^{T}U=UU^{T}=I$, and $D$ is a diagonal matrix containing the eigenvalues $\lambda_{i}$ of $\mathcal{M}$.
-A normalized vector $\mathbf{c}$ can be expressed by the eigenvectors $\mathbf{e}_{i}$: $\mathbf{c}=\sum_{i}n_{i}\mathbf{e}_{i}.$ Substituting this into the Rayleigh quotient $W$, we obtain:
-$W(\mathcal{M},\mathbf{c})=$ $\sum_{ij}n_{i}n_{j}\mathbf{e}_{i}^{T}UDU^{T}\mathbf{e}_{j}$ (3)
-$=$ $\sum_{ij}n_{i}n_{j}D_{ij}=\sum_{i}n_{i}^{2}\lambda_{i}.$
+A normalized vector $\mathbf{c}$ can be expressed by the eigenvectors of $\mathcal{M}$: $\mathbf{c}=\sum_{i}n_{i}\mathbf{e}_{i}.$ Substituting this into the Rayleigh quotient $W$, we obtain:
+$W(\mathcal{M},\mathbf{c})=\sum_{ij}n_{i}n_{j}D_{ij}=\sum_{i}n_{i}^{2}\lambda_{i}.$ (3)
 This indicates that if there is a subset of eigenvectors $Q$, where $\forall\mathbf{e}_{i}\in Q$, $W(\mathcal{M},\mathbf{e}_{i})>0$, then any linear combination $\mathbf{c}=\sum_{i}n_{i}\mathbf{e}_{i}$ will also satisfy $W(\mathcal{M},\mathbf{c})>0$.
@@
 Step 3.—Find the suitable operator to witness entanglement within the intersection space.
-The operators for detecting entanglement structure can be optimized by maximizing the minimum value of $W(\mathcal{M}_{i},\mathbf{c})$ across all matrices $\mathcal{M}_{i}$,
+The operators for detecting entanglement structure can be optimized by maximizing the minimum value of $W(\mathcal{M}_{l},\mathbf{c})$ across all matrices $\mathcal{M}_{l}$,
 $g=\min\{W(\mathcal{M}_{1},\mathbf{c}),W(\mathcal{M}_{2},\mathbf{c}),W(\mathcal{M}_{3},\mathbf{c}),\ldots\},$ (4)
-with $\mathcal{M}_{i}$ are associated to each $\mathcal{K}$-partition.
-However, this optimization can be analytically challenging with the existence of different Rayleigh quotients $W(\mathcal{M}_{i},\mathbf{c})$.
-To simplify the optimization, we construct a combined function that integrates different $\mathcal{M}_{i}$
-$f=$ $\sum_{i}W(\mathcal{M}_{i},\mathbf{c})=W\left(\sum_{i}\mathcal{M}_{i},\mathbf{c}\right)$ (5)
+with $\mathcal{M}_{l}$ are associated to each $\mathcal{K}$-partition.
+However, this optimization can be analytically challenging with the existence of different Rayleigh quotients $W(\mathcal{M}_{l},\mathbf{c})$.
+To simplify the optimization, we construct a combined function that sums over different $\mathcal{M}_{l}$.
+$f=$ $\sum_{l}W(\mathcal{M}_{l},\mathbf{c})=W\left(\sum_{l}\mathcal{M}_{l},\mathbf{c}\right)$ (5)
 $=$ $\sum_{i=1}^{k}\sum_{j=1}^{k}n_{i}n_{j}(\mathbf{p}_{i}^{T}M\mathbf{p}_{j})=\mathbf{n}^{T}\mathcal{Q}\mathbf{n}\;,$
@@
 Here, $\mathcal{Q}$ is a symmetric matrix with elements $q_{ij}=\mathbf{p}_{i}^{T}\mathcal{M}\mathbf{p}_{j}$.
-The optimal $\mathbf{c}$ can be found by maximizing the value of $g$ in Eq.
-(4).
-This involves selecting the positive eigenvalues $\lambda_{j}$ of $\mathcal{Q}$ that satisfy $g>0$ and then determining the eigenvector corresponding to the optimal $\lambda_{j}$ that maximizes $g$.
-This ends Step 3.
+To determine the maximum value of $g$ in Eq.
+(4) from the vector $\mathbf{c}$, we select all positive eigenvalues $\lambda_{j}$ of $\mathcal{Q}$ that satisfy $g>0$.
+For each of these positive eigenvalues, we examine their corresponding eigenvectors and determine the optimal eigenvector $\mathbf{n}$ that yields the maximum value of $g$.
+This concludes Step 3.
 These three steps provide a general method for finding an operator to witness entanglement for different partitions.
-In the following, we will apply this methodology to various general CV states (both Gaussian and non-Gaussian), demonstrating its effectiveness in characterizing multipartite entanglement structures.
+In the following, we apply this methodology to various general CV states (both Gaussian and non-Gaussian), demonstrating its effectiveness in characterizing multipartite entanglement structures.
 Figure 2: Multipartite entanglement structure detection.
@@
 This procedure results in states that includes the most common Gaussian and non-Gaussian states prepared in experiments, such as squeezed and photon subtracted states (see details in the supplementary material [51]).
-The entanglement structure of these random generated multimode states cannot be characterized by the previous criteria based on covariance matrix [38, 39, 40, 36, 37] and QFI [43, 47].
+The entanglement structure of these random generated multimode states cannot be characterized by the previous criteria based on the covariance matrix [37, 38, 39, 35, 36] and QFI [45, 46].
 The primary limitations of these methods are their inability to detect entanglement across different partitions due to their focus on a specific partition, or their reliance on measuring variances of linear combinations of position and momentum operators, which do not capture non-Gaussian correlations.
 In contrast, our method based on QFI overcomes these challenges.
-To test the effectiveness of our method, We randomly generate $40000$ four-mode quantum states (ensuring 10,000 states for each type of entanglement structure).
+To test the effectiveness of our method, We randomly generate $40,000$ four-mode quantum states (ensuring 10,000 states for each type of entanglement structure by Peres–Horodecki criterion [56, 57]).
 As shown in Fig.
@@
 For four modes, using first-order encoding operators $\hat{A}$ (combinations of $\hat{x}$ and $\hat{p}$) we can reach over 90% detection rate.
-By including higher-order operators ($\hat{x}^{2},\hat{p}^{2},\hat{x}\hat{p}+\hat{p}\hat{x},\dots$), the successful rate of characterizing entanglement structure increases to almost 100%.
+Moreover, by including higher-order operators ($\hat{x}^{2},\hat{p}^{2},\hat{x}\hat{p}+\hat{p}\hat{x},\dots$), the successful rate of characterizing entanglement structure increases to almost 100%.
 Figure 3: Witnessing fully inseparable entanglement for different non-Gaussian states.
@@
 Intriguingly, we demonstrate that extending the analysis to include second-order operators significantly enhances the successful detection rate, showing robustness against channel loss.
-This improvement is depicted in the green curves of Fig.
+This improvement is depicted in the solid green curves of Fig.
 3(b).
 In addition to examining non-Gaussian states generated from core states, we also explore non-Gaussian states produced by nonlinear processes, such as the three-photon SPDC process.
-This process has been realized experimentally [27, 26] and has theoretically attracted significant interest due to its challenges in characterizing its correlations [57, 58, 59, 60, 61, 62].
-For our analysis, we randomly generated $10^{4}$ non-Gaussian states using the three-photon SPDC process, described by the Hamiltonian $\hat{H}=(\chi_{1}\hat{a}\hat{b}^{2}+\chi_{2}\hat{b}\hat{c}^{2}+\chi_{3}\hat{c}\hat{a}^{2})+h.c.$ Here the parameters $\chi_{i}t$ are randomly selected from the interval [0, 0.04], and the initial states are vacuum states.
+This process has been realized experimentally [26, 25] and has theoretically attracted significant interest due to its challenges in characterizing its correlations [59, 60, 61, 62, 63, 64, 65].
+Our method is not only applicable to these specific SPDC processes, but also more broadly.
+To demonstrate it, we randomly generated $10^{4}$ non-Gaussian states using the three-photon SPDC process, described by the Hamiltonian $\hat{H}=(\chi_{1}\hat{a}\hat{b}^{2}+\chi_{2}\hat{b}\hat{c}^{2}+\chi_{3}\hat{c}\hat{a}^{2})+h.c.$ Here the parameters $\chi_{i}t$ are randomly selected from the interval [0, 0.04], and the initial states are vacuum states.
 As illustrated in Fig.
@@
 In contrast, higher-order QFI successfully identifies entanglement.
+The reason the detection rate of fully inseparable states does not reach 100% is that the randomly selected $\chi_{i}$ values have some probability of being very small, resulting in states with no or very weak multimode correlations.
 These results exhibit the versatility of our method across different non-Gaussian states, demonstrating its effectiveness in accurately identifying operators that capture complex non-Gaussian correlations.
-In conclusion, we provide a general method for characterizing entanglement for arbitrarily Gaussian and non-Gaussian states.
-Based on QFI, we establish a criterion and propose a systematic approach to identify entanglement by analytically optimizing operators from accessible family of operators.
-This method involves three key steps: (i) identifying convex sets of operators that can be used to witness $\mathcal{K}$-partition entanglement, (ii) determining the intersection of these sets, and (iii) selecting a suitable operator at the intersection.
+Conclusion.— We provided a general method for characterizing entanglement for arbitrarily Gaussian and non-Gaussian states.
+Based on the QFI, we established a systematic approach to analytically identify an operator to efficiently capture different quantum correlations.
+This method involves three key steps: (i) identifying convex sets of operators that can be used to witness $\mathcal{K}$-partition entanglement, (ii) determining the intersection of these sets, and (iii) selecting an optimal operator at the intersection.
 To verify the accessibility and effectiveness of our method, we randomly generated $10^{5}$ states based on core states and nonlinear processes, which demonstrate a very high proportion of success in detecting entanglement structure.
 Furthermore, we exhibit the advantage of our method in resisting channel loss and detecting non-Gaussian entanglement.
-This approach provides a systematic way to distinguish entanglement structure for arbitrary states in CV systems, paving the way for further exploration and detection of entanglement in multimode experimental systems.
+This approach provides a systematic way to distinguish entanglement structure for arbitrary states in CV systems.
+Additionally, our approach requires only a single encoding operator in QFI to capture various correlations, making it more practical for characterizing entanglement structures in experiments.
+Recent work has successfully applied our method to characterize multimode entanglement using only homodyne measurements [66].
+Furthermore, applying the optimal operator to the nonlinear squeezing parameter [67] offers a potential experimental pathway for detecting entanglement structures.
+Our findings open new avenues for exploring and detecting entanglement in multimode experimental systems.
 Acknowledgements.
@@
 208886, and by The Branco Weiss Fellowship – Society in Science, administered by the ETH Zürich.
+This work was supported by the project PID2023-152724NA-I00, with funding from MCIU/AEI/10.13039/501100011033 and FSE+, by the project RYC2021-031094-I, with funding from MCIN/AEI/10.13039/501100011033 and the European Union ‘NextGenerationEU’ PRTR fund, by the Ministry of Economic Affairs and Digital Transformation of the Spanish Government through the QUANTUM ENIA Project call—QUANTUM SPAIN Project, by the European Union through the Recovery, Transformation and Resilience Plan—NextGenerationEU within the framework of the Digital Spain 2026 Agenda, and by the CSIC Interdisciplinary Thematic Platform (PTI+) on Quantum Technologies (PTI-QTEP+).
+This work was supported through the project CEX2023-001292-S funded by MCIU/AEI.